Drew owns and operates an onion packing plant. To help reduce costs on labor and to increase efficiency, Drew is considering purchasing an automatic 50 lb. bagging machine and a machine that automatically stacks the 50 lb. bags onto pallets.   The cost of the two machines combined and their installation is $300,000. Drew will make a 20% down payment and finance the rest of the machinery with an amortized loan over 15 years at a 5.5% interest rate.   Drew predicts that by using these two machines, his plant will be able to increase output, therefore increasing operating receipts by $30,000 per year.   Also, this machinery will save him approximately $10,000 in labor costs each year by the increase in efficiency.   However, other operating expenses such as electricity and insurance will increase by approximately $5,000 per year.   Drew assumed a straight-line depreciation over 15 years and the life of the investment is 7 years.   The terminal value is $180,000 after the 7 years, and Drew requires a pretax 10% rate of return to capital.   The marginal tax rate over the next 10 years is 25%.

(i)           What is the initial cost for the machine?

a.$60,000                                               b.  $300,000

c.$240,000                                            d.$150,000

Enter response here

(ii)         What are the after-tax net returns?

a.$22,500                                               b.$35,000

c.$26,250                                            d.$15,000

Enter response here

(iii)        What are the tax savings from depreciation?

a.$0                                                         b.$10,000

c.$15,000                                               d.$5,000

Enter response here

(iv)        What is the after-tax terminal value?

a.$180,000                                          b.$300,000

c.$160,000                                          d.$175,000

Enter response here

(v)         What is the appropriate discount rate to calculate the NPV given the above information?

a.10%                          b.25%

c.5.5%                         d.7.5%

Enter response here

(vi)        What is the net present value?

a.$23,910.14                                     b.($28,999.10)

c.$28,999.10                                       d.$26,250

Enter response here

(vii)What is the least additional operating receipts for the new machine.

a.   $37,273                         b.    $32,089

c.    $47,506                           d.    $45,695  

Enter response here
Answer #1

Answer

i. Initial cost of the machine is $300000*20% ( Down payment)=$60,000

Option a is correct

ii .After tax net return

Net operating income=30000

Additional labors savings=10000

Total savings =40000

less Expenses=5000

= Net profit=40000-5000=35000

less tax @ 25%= 8750

After Tax net return =26250.

Option c

iii.Option d 5000

Depreciation= 300000/15=20000

Tax on depreciation=20000*25%=5000

iv.option c

depreciation per year=20000

for 7 years 7*20000=140000

Total value -depreciation

300000-140000=160000

v. Option d 7.5%

Rate of interest - tax shield

10-25%=7.5%